OpenStudy (anonymous):

Solve: -2x - 13 = 8x + 7 -2 2 -6 6

OpenStudy (anonymous):

-2x - 8x = 13 + 7

OpenStudy (ivettef365):

combine like terms, when switch from one side to another change signs as follows: -13 -7 = 8x + 2x ( do you see what I did ? )

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

-2 x - 8x = ? 13+7 = ?

OpenStudy (samigupta8):

20=10 x u will get aftr dat...

OpenStudy (anonymous):

then you will find the answer

OpenStudy (samigupta8):

sry dat was 20=-10x

OpenStudy (anonymous):

-2x - 13 = 8x + 7 I would say Add 13 both sides -2x - 13 = 8x + 7 13 13 ========== -2x=8x=20 Now bring 8x left side -2x-8x=20 -10x=20 Divide 20 with -10 x=20/-10 x=-2

OpenStudy (samigupta8):

x=-2

OpenStudy (anonymous):

@deezgurls

OpenStudy (anonymous):

I got -2 from @PFEH.1999 instructions but @Munch. you explained it clearly too

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

your welcome :)